# MAViS: A Multi-Agent Framework for Long-Sequence Video Storytelling

### v4

@@ Appendix C Implementation Details
All agents in our framework are implemented using AutoGen Wu et al. (2024) and powered by GPT-4o Achiam et al. (2023). The T2I model pool comprises FLUX.1 BlackForestLabs (2023), Stable Diffusion 3.5 Rombach et al. (2022), and GPT-Image OpenAI (2025), while the I2V model pool includes Veo2 Google (2024), Gen-3 Runway (2024a), Gen-4 Runway (2024b), LUMA Luma (2024), HunyuanVideo-I2V Kong et al. (2024), and Wan2.1 Ang and Baole (2025). We use MUDI Jang et al. (2024) for LoRA training and Hailuo MiniMax (2024) for the T2A generation. The maximum number of 3E iterations is set to 4 for script writing, shot designing, and voice designing; 2 for the T2I process; 1 for the I2V process; and 5 for subtitle refining. As the first work targeting the efficient translation of single lines of ideas to long-sequence video storytelling, no suitable benchmark dataset exists. Therefore, we constructed a test set of 20 user prompts, each targeting a one-minute video for evaluation.
@@
The T2I model pool includes the open-source models FLUX BlackForestLabs (2023) and Stable Diffusion 3.5 StabilityAI (2024), as well as the API-based GPT-Image OpenAI (2025). The I2V model pool comprises open-source models Hunyuan Kong et al. (2024) and Wan2.1 Ang and Baole (2025), along with API-based models Gen3 Runway (2024a), Gen4 Runway (2024b), LUMA Luma (2024), and Veo2 Google (2024). For T2A, we use the Minimax Hailuo API MiniMax (2024). All experiments were conducted on one NVIDIA V100 GPUs.
@@ C.2 Baselines
We compare MAViS with three long video generation baselines: VGoT, Mora, and MovieAgent. Among them, Mora and MovieAgent adopt a modular design and are equipped with the same image and video generation models as MAViS, and are also powered by GPT-4o. Since Mora and MovieAgent do not support script generation, we provide them with scripts formed by our script writing stge. For MovieAgent, which requires manual training of LoRA models, we generate character images based on the provided scripts and train LoRA models for each script.
@@ Appendix D Compare with More Baselines
We first compare our method with DreamFactory Xie et al. (2024). Given the scarcity of real videos in long-sequence video storytelling with assistive generation, computing the FID metric is not feasible. As shown in Table 7, MAViS achieves higher scores across most metrics; however, there is a substantial gap in Inception Score. We adopted a widely used implementation of the Inception Score, which may differ from the version used by DreamFactory, potentially leading to differences in score range. Since the baseline is not open-sourced, we are unable to verify the exact cause of this discrepancy.
@@ Appendix E User Study
In the User Study, participants were asked to vote for the best video in each group based on seven evaluation criteria. As listed in Table 10, the criteria include: Narrative Expressiveness, Visual Quality, User Prompt Alignment, Character Consistency, Character Naturalness, Background Consistency, and Background Realism. Prior to voting, detailed guidelines for each metric were provided to ensure consistency and reliability in evaluation. Some screenshots of the user study HTML interface is shown in Figures 3, 4, and 5.
@@ Appendix F Evaluation on the Existing Dataset
We are establishing an efficient agentic system from users’ single line of ideas to final pixels plus other multimodalities. However, existing long video benchmarks, MovieBench Wu et al. (2025a) and ViStoryBench Zhuang et al. (2025), focus only on benchmarking T2I and I2V models’, not on the entire agentic system. In particular, they ignore the script generation and scripts’ compatibility with existing T2I and I2V models’ capabilities.
@@ Appendix G Quality-Efficiency Analysis
For script writing, as shown in Tabel 12, the performance is already satisfactory at . Increasing further can improve generation quality with minimal impact on overall efficiency.
@@
For keyframe generation, as shown in Tabel 13, the performance is already satisfactory at . Increasing further can enhance generation quality, with only a slight impact on overall efficiency.
For video animation, as shown in Tabel 14, the performance is already satisfactory at . While increasing can further improve quality, it has a significant impact on overall efficiency. Therefore, we set for video animation in our experiments.
For shot designing, voice designing, and subtitle refining, as shown in Tabel 15, increasing improves generation quality with minimal impact on overall efficiency. However, since errors in these stages can critically affect the entire generation process, we set for these tasks in our experiments.

### v5

@@ Appendix C Implementation Details
All agents in our framework are implemented using AutoGen Wu et al. (2024) and powered by GPT-4o Achiam et al. (2023). The T2I model pool comprises FLUX.1 BlackForestLabs (2023), Stable Diffusion 3.5 Rombach et al. (2022), and GPT-Image OpenAI (2025), while the I2V model pool includes Veo2 Google (2024), Gen-3 Runway (2024a), Gen-4 Runway (2024b), LUMA Luma (2024), HunyuanVideo-I2V Kong et al. (2024), and Wan2.1 Ang and Baole (2025). We use MUDI Jang et al. (2024) for LoRA training and Hailuo MiniMax (2024) for the T2A generation. The maximum number of 3E iterations is set to 4 for script writing, shot designing, and voice designing; 2 for the T2I process; 1 for the I2V process; and 5 for subtitle refining. As the first work targeting end-to-end long-sequence video storytelling, no suitable benchmark dataset exists. Therefore, we constructed a test set of 20 user prompts, each targeting a one-minute video for evaluation.
@@
The T2I model pool includes the open-source models FLUX BlackForestLabs (2023) and Stable Diffusion 3.5 StabilityAI (2024), as well as the API-based GPT-Image OpenAI (2025). The I2V model pool comprises open-source models Hunyuan Kong et al. (2024) and Wan2.1 Ang and Baole (2025), along with API-based models Gen3 Runway (2024a), Gen4 Runway (2024b), LUMA Luma (2024), and Veo2 Google (2024). For T2A, we use the Minimax Hailuo API MiniMax (2024). All experiments were conducted on two NVIDIA H100 GPU.
@@ C.1 Metrics
Compliance Rate
Compliance Rate is an objective and transparent metric defined as the fraction of outputs that are fully compliant with the guidelines (Line 596). A “fully compliant” output is defined as follows: Shot Designing: All seven fields are present in the correct format. Voice Designing: Downloadable BGM consistent voice ID correct format. Subtitle Refining: Narration duration video duration.
If any of the above rules are violated, the workflow will produce an error. Therefore, the Compliance Rate directly measures the necessity and effectiveness of the 3E principle across these three stages, and there is no better metric that can faithfully reflect this property.
@@ C.2 Baselines
We compare MAViS with three long video generation baselines: VGoT, Mora, and MovieAgent. Among them, Mora and MovieAgent adopt a modular design and are equipped with the same image and video generation models as MAViS, and are also powered by GPT-4o. Since Mora and MovieAgent do not support script generation, we provide them with scripts produced by our script writing stage. For MovieAgent, which requires manual training of LoRA models, we generate character images based on the provided scripts and train LoRA models for each script.
@@ Appendix D Compare with More Baselines
We first compare our method with DreamFactory Xie et al. (2024). Due to the lack of real videos in automatic long-sequence video storytelling, it is not feasible to compute the FID metric. As shown in Table 8, MAViS achieves higher scores across most metrics; however, there is a substantial gap in Inception Score. We adopted a widely used implementation of the Inception Score, which may differ from the version used by DreamFactory, potentially leading to differences in score range. Since the baseline is not open-sourced, we are unable to verify the exact cause of this discrepancy.
@@ Appendix E User Study
In the User Study, participants were asked to vote for the best video in each group based on seven evaluation criteria. As listed in Table 11, the criteria include: Narrative Expressiveness, Visual Quality, User Prompt Alignment, Character Consistency, Character Naturalness, Background Consistency, and Background Realism. Prior to voting, detailed guidelines for each metric were provided to ensure consistency and reliability in evaluation. Some screenshots of the user study HTML interface is shown in Figures 6, 7, and 8.
@@ Appendix F Evaluation on the Existing Dataset
We are establishing an end-to-end agentic system from users’ single line of story theme to final pixels plus other multimodalities. However, existing long video benchmarks, MovieBench Wu et al. (2025a) and ViStoryBench Zhuang et al. (2025), focus only on benchmarking T2I and I2V models’, not on the entire agentic system. In particular, they ignore the script generation and scripts’ compatibility with existing T2I and I2V models’ capabilities.
@@ Appendix G Quality-Efficiency Analysis
For script writing, as shown in Table 13, the performance is already satisfactory at . Increasing further can improve generation quality with minimal impact on overall efficiency.
@@
For keyframe generation, as shown in Table 14, the performance is already satisfactory at . Increasing further can enhance generation quality, with only a slight impact on overall efficiency.
For video animation, as shown in Table 15, the performance is already satisfactory at . While increasing can further improve quality, it has a significant impact on overall efficiency. Therefore, we set for video animation in our experiments.
For shot designing, voice designing, and subtitle refining, as shown in Table 16, increasing improves generation quality with minimal impact on overall efficiency. However, since errors in these stages can critically affect the entire generation process, we set for these tasks in our experiments.
@@
| Model/API | Cost | Time overhead | Duration | GPU |
| Veo 2 | 2.8$ | 31.7s | 8s | - |
| Gen 3 | 0.5$ | 23.7s | 10s | - |
| Gen 4 | 0.5$ | 40.8s | 10s | - |
| Luma | 1.27$ | 36.2s | 5s | - |
| Hunyuan-I2V | - | 28.7min | 5s | 1 H100 |
| Wan 2.1 | - | 36.3min | 3s | 1 H100 |
Appendix H Generation Time Overhead.
Using all models in the model pool, the average time for MAViS to generate a long video on two H100 GPUs is 13.63 hours. In contrast, it will take our internal artists 5 days to complete a pipeline similar to ours: data generation, evaluation, selection, and composition. The process can be significantly accelerated if only API-based I2V models are used. The resources required for a single video model to generate one 720p shot are as shown in Table 17.
Appendix I Most Common Issues and Success Rate
We report the most common issues and success rates across all the stages that apply the 3E principle during generation:
Script writing: Adjacent shots are set in the same or strongly connected locations, violating the script writing guidelines.
Shot designing: Incomplete content, such as insufficient background detail.
T2I: Facial and limb distortions.
I2V: Abnormal character movements.
Voice designing: BGM download failures.
Subtitle refinement: The voice duration exceeds the clip length.
Using criteria such as avoiding visual tearing, distortion, physically implausible motion, ID inconsistency, and style inconsistency as overall success measures, MAViS achieves 0.8621, compared to 0.4875 without the 3E principle.